OpenStudy (anonymous):

write the equation for the exponential function that passes through the points (0.5) and (1,15)

OpenStudy (shadowfiend):

Hm. I assume an exponential function means in the form \(y = ae^x\)?

OpenStudy (anonymous):

yes

OpenStudy (shadowfiend):

Ok, so you basically have two equations and two unknowns if you plug in your two values: \[\begin{align} y &= ae^x\\ 5 &= ae^0\\ 15 &= ae^1 \end{align}\] You have a huge advantage because one of the equations is \(ae^0\). What's e^0?

OpenStudy (anonymous):

hey shadow

OpenStudy (anonymous):

shadow youre missing something. y = Ae^(kt). there should be an A and a k. or two parameters. Shadow, you have one parameter , just the a.

OpenStudy (anonymous):

or if you want to use x, then y = Ae^(kx)

OpenStudy (shadowfiend):

Absolutely right -- had looked at it a little further earlier and there were some holes in the reasoning. teekay hadn't gotten back yet, though, so I figured I'd wait. Thanks!